OpenStudy (anonymous):

How would you simplifiy 1/4sqrt5 ?

OpenStudy (anonymous):

I think my notation is unclear

jhonyy9 (jhonyy9):

sqrt5/20

OpenStudy (saifoo.khan):

\[\frac{1}{4 \sqrt5 } \times \frac{4 \sqrt5}{4 \sqrt5}\]

OpenStudy (anonymous):

1/(4sqrt5)

jhonyy9 (jhonyy9):

,,saifoo.khan" really sure ???

OpenStudy (anonymous):

jhonyy9, i checked the answer, you are correct. thanks

jhonyy9 (jhonyy9):

ok

OpenStudy (saifoo.khan):

isnt it going to be solved that way?

jhonyy9 (jhonyy9):

this need to simplify

OpenStudy (saifoo.khan):

Yeah, but i showed the steps since u already showed the answer! :)

jhonyy9 (jhonyy9):

ok

OpenStudy (saifoo.khan):

Hmm.